# FLex&Chill: Improving Local Federated Learning Training with Logit Chilling

### v1

@@ 3.2 FLex&\&&Chill and Logit Chilling
4 Experiment
We evaluate the performance of FLex&\&&Chill using three datasets: (i) FEMNIST from the LEAF database Caldas et al. (2018b), including the Extended MNIST handwriting data Cohen et al. (2017) , (ii) the CIFAR10 dataset, and (iii) the CIFAR100 dataset Krizhevsky et al. (2009). As metrics, we use the average accuracy and training loss observed at the clients, selected to observe both the performance gain and the effectiveness of FLex&\&&Chill in its training operations. Furthermore, we also present the number of federated learning rounds needed to achieve a pre-defined target accuracy.
(a) FEMNIST-DNN
(b) CIFAR10-CNN
(c) CIFAR100-ResNet18
4.1 Experiment Setup
Dataset. We detail the datasets used in our work as follows.
  • FEMNIST dataset: Sourced from the LEAF dataset, the FEMNIST dataset comprises of 10 handwritten digits (0-9) and 52 characters (26 lowercase and 26 upper case) images contributed by 712 users, totaling 157,132 samples. We selected a subset of 36 users, each with diverse amounts of local data to account for the non-iid environment. We followed the implementations provided by the original work Caldas et al. (2018a).
  • CIFAR10 and CIFAR100 dataset: We use the baseline CIFAR10 and CIFAR100 datasets Krizhevsky et al. (2009) and manually partition the data over 50 federated learning users, with each user holding 1,000 samples.
Experiment Details. Each case underwent local training for 300 federated training rounds, utilizing FedAvg McMahan et al. (2017) as the baseline federated learning framework. FedAvg is a widely used framework for federated learning. In each round, the server selects 10 participating clients, employing Stochastic Gradient Descent (SGD) as the optimizer with a learning rate of 0.001 Ruder (2016). Each client trained its model for 10 local epochs, employing a batch size of 16, and utilized cross-entropy loss as the loss function.
4.2 Average Client-side Accuracy
(a) FEMNIST-DNN
(b) CIFAR10-CNN
(c) CIFAR100-ResNet18
Figure 5: Federated learning client-side model training loss observed for different training temperature values. We can notice that overall a lower training loss is observed when lower temperature values are used for model training.
We report the average client-side model accuracy results obtained for models trained with different temperatures in Figure 4. For all three datasets (and the three respective models used for each test case) we see an increasing accuracy trend observed for lower temperatures. Note that for the FEMNIST-DNN case with $T$=0.05, we noticed that the accuracy converges after 120 rounds; thus, perform an early exit for this case. One interesting aspect we can notice is that the accuracy for the $T$=0.05 case, which is the lowest temperature we test for, does not show the best performance for the CIFAR100-ResNet18 experiment. While the $T$=0.05 case shows the best performance of FEMNIST-DNN and CIFAR10-CNN, for CIFAR100-ResNet18, we see a accuracy worse than the $T$=0.5 case. This suggests that training with the lowest possible temperature value is not always ideal, rather this result highlights the importance of identifying a proper training temperature as we will further discuss in Section 5. A closer zoom into the plots suggest that the best performing configuration can change over the training round and can vary for different datasets, but overall, we can notice that exploiting a temperature value lower than “1” in the training process can help improve the classification accuracy.
The average accuracy for all federated learning clients after 300 training rounds presented in Table 1 confirm our observations, indicating that with $T$=0.05 the CIFAR10 case shows a 3.37% improvement in accuracy compared to the $T$=1 case. Even for other cases, the use of a low temperature value show promising results in terms of accuracy. We emphasize once more that this does not mean that the accuracy will always perform better at the lowest temperature. Rather, our results suggest the need to carefully explore the fractional temperature space in the federated learning training process.
4.3 Training Loss
Next we compare the training loss observed at the clients. In Figure 5 we present the average the loss observed at each federated learning training round. Again, we note that we perform early exit for the FEMNIST-DNN case at 120 rounds when $T$=0.05 due to quick model convergence. We can notice there that the use of lower temperatures in the training phase generally helps quickly lower the training loss during the local training phase for federated learning clients. At the same time, when observing the CIFAR100-ResNet18 plots, we can see that the loss for $T$=0.05 does not converge well. In fact, this serves as a reason behind the lower accuracy reported for CIFAR100-ResNet18.
From the quantitative values in Table 1 where we plot the average loss observed at all federated learning clients after 300 training rounds (120 for the FEMNIST-DNN @ $T$=0.05), we can make similar observations. Especially, for FEMNIST-DNN, we see a six-fold improvement in loss values compared to $T$=1. For the CIFAR100-ResNet18 case, however, we notice that the lowest loss convergence is in fact the $T$=0.25 case rather than 0.05. This observation confirms the need to properly consider and explore fractional temperature values for client-side model training.
4.4 Federated Learning Convergence Analysis
Table 1: Average classification accuracy, model training loss, and number of federated learning training rounds needed to achieve the maximum accuracy of the $T$=4 case for each dataset-model combination. Quantitative results suggest the effectiveness of exploiting low temperatures for model training in federated learning both in terms of model accuracy and training efficiency.
Centralized - iid Federated - iid Federated - non-iid $T$ Acc. (%) Loss Acc. (%) Loss Acc. (%) Loss 4.0 58.05 1.152 48.22 1.473 37.58 1.724 2.0 63.27 1.041 51.90 1.403 40.35 1.638 1.0 65.18 1.044 54.66 1.325 41.34 1.594 0.5 64.19 1.030 53.19 1.320 41.52 1.551 0.25 63.88 1.092 54.64 1.265 43.41 1.532 0.05 60.98 1.227 47.60 1.701 44.71 1.528
Table 2: Average inference accuracy and model training loss for CIFAR10-CNN with different configurations. The impact of FLex&\&&Chill is prominent for federated learning scenarios with non-iid datasets, whereas its impact is less when elsewise.
Finally, the last columns for each test case in Table 1 we present the number of federated learning training rounds needed to achieve a preset target accuracy for different temperature values using for model training. This number provides an understanding of how efficient the local training processes are when different temperature values are applied. S
[…]
From the results in Table 1 we can notice that with $T$=0.05 the FEMNIST-DNN case achieves the target accuracy after only 22 training rounds. We see similar trends of low temperatures positively affecting the training efficiency throughout our results. This is from the generally aggressiveness of training behaviors when applying lower temperature training. Given a single training sample, while high temperature-based training conservatively absorbs the available information, lower temperatures are more progressive in making model changes. Given the small datasets used for local training and the need to effectively deliver its knowledge to the server for aggregation, we see FLex&\&&Chill as an effective way to train client-side models in the federated learning process.
We now examine the impact of FLex&\&&Chill on different model training configurations. Specifically, using the CIFAR10-CNN test case, we examine two additional configurations: (i) centralized learning with iid data and (ii) federated learning with iid data. Note that the federated learning network is identical to our previous experiments except for the dataset distribution and was tested for 300 rounds. By comparing these results with the non-iid federated learning results, we target to understand the impact of FLex&\&&Chill on federated learning and dataset disparity in greater detail.
As Table 2 shows, applying FLex&\&&Chill, thus a lower $T$ does not induce a positive impact on the performance of centralized learning scenarios. In fact, for centralized learning, controlling $T$ for the training process does not show a noticeable impact. We see the same phenomena for the federated learning system with iid samples as well, suggesting that the performance enhancements we see in Table 1 is not an effect of the federated learning process itself. We do point out that holding iid datasets in federated learning will show overall higher performance compared to non-iid configurations (regardless from $T$), but is not a practical case to assume when deploying federated learning systems. Rather, as we discussed, the impact of exploiting low training temperatures is beneficial when dealing with non-iid datasets.
Figure 6: Accuracy drop observed before and after server-side model aggregation. The aggregation process of client models trained with lower temperatures better retain the individual knowledge.
4.6 Model stability of low temperature training during federated learning aggregation
As discussed in Section 3, low temperatures can expedite local training even when dealing with a small amount of training data. Nevertheless, the effect of low temperature training on the model generalization and aggregation stability of federated learning has not yet been thoroughly studied. To inspect the stability of federated learning aggregation for models trained with low temperatures, we measured the the performance change of each client-side model before and after the server-side aggregation round.
Figure 6 plots the changes in accuracy observed for both participating and non-participating clients at each federated learning round. Specifically, once the model is aggregated, each client computes its accuracy with both the pre-update and the post-update model and computes their difference. Note that pre-update models for clients participating in the round would have updated their local models (very recently) with their own data. Thus, the pre-model performance will be high. However, since each model is trained aggressively over heterogeneous data, the aggregation process could puddle up this knowledge and the aggregated model may lose information. Ideally, with their local information properly aggregated, the performance drop will be minimal. Non-participating clients on the other hand, would hope to have an improved updated model the federated learning process.
As Figure 6 shows the models trained with low temperatures retain the knowledge better than the medium and hot models for participating clients (solid plots groups at bottom). At the early rounds, the models with high temperature show less performance drop, but the fast learning speed of low temperature model amortizes the performance drop, which results in faster learning of the aggregated model. Furthermore, we can also notice that the aggregation process does not interfere with the performance of non-participating clients (grouped on top), implying that federated learning with low temperature training preserves the generalization power.
5 Discussion
$\bullet$ Identifying the optimal temperature. Our study shows the effectiveness of exploiting lower temperatures as a way to enhance federated learning performance. However, an important relevant question remains unanswered: “How do we determine the appropriate or optimal temperature for different applications?” With varying dataset and models that different federated learning application use, the ideal temperature value will also differ. The scope of this work is on identifying and quantifying the potential benefits of lower temperature-based local training in federated learning and we leave the task of identifying per-purpose optimal temperatures to individual system designers and future research.
$\bullet$ Dynamic per-sample temperature scaling. While our paper focuses on understanding the advantages of maintaining a constant low temperature for the training process, this may not be the ideal approach when targeting maximum performance. In reality, varying temperatures can be applied in the training process with respect to their significance Shin et al. (2022); Zeng et al. (2021). That would mean that the training process can exploit lower temperatures for the samples that is “harder” to learn, while higher temperatures are used for those easier to understand. Alternatively, a more course-grain approach of dynamically adjusting the temperature on a per-federated learning round-basis can be another meaningful approach to adapt towards changing local data characteristics.
$\bullet$ Supporting heterogeneous federated learning. While many federated learning frameworks assume homogeneous devices, recent work have pointed out that this is not always the case Park and Ko (2023). In a federated learning network consisting of devices with heterogeneous computing/network capabilities, the issue of “stragglers”, which are devices that take longer to complete local training, becomes important to address. Leveraging insights gleaned from our empirical observations, applying lower temperatures for potential stranglers can be an interesting way of uniforming the local training duration. This would require a prior analysis on the conditions expected at each device and a global consensus on the temperature configuration. We find this as an interesting topic that requires further investigation.
6 Conclusion
This research introduces a novel training approach for federated learning systems, FLex&\&&Chill, aimed at enhancing convergence and improving model performance by exploiting a temperature scaling method termed logit chilling in federated learning local model training operations. Specifically, our work revolves around the implications of employing fractional temperature values ($10<T<$ 1) during the model training process. Our findings, derived from an extensive set of evaluations using three datasets and three widely applied neural networks for federated learning, demonstrate that exploiting the FLex&\&&Chill approach in the client-side model training process holds the potential to expedite model convergence by inducing more pronounced changes despite a limited number of local samples. Consequently, we show that by doing so, we can also achieve an overall higher inference accuracy. While the impact of lower temperature vary with respect to the neural network architecture and dataset, our evaluations reveal up to 6.00$\times$ improvement in federated learning model convergence speed and 3.37% improvement in inference accuracy. We see this work as a first step into devising a more fine-tuned (and optimal) training mechanism for federated learning systems.

### v2

@@ C.1 FLex&\&&Chill As An Algorithm
C.2 FLex&\&&Chill with Advanced FL Schemes
As mentioned in Section 3.2, FLex&\&&Chill is an orthogonal approach to enhance the federated learning process. To demonstrate this, we trained the 2-layer CNN with FedProx Li et al. [2020] and SCAFFOLD Karimireddy et al. [2020] representing more recently proposed FL schemes compared to FedAvg used in this work.
Hyperparameters FedProx SCAFFOLD total clients 50 communication rounds 300 optimizer SGD learning rate 0.001 local epoch 5 participants per round 5 batch size 32 $\alpha$ 0.5 loss function Cross Entropy $\mu$ 0.01 N/A
Table 6: Default hyperparameters used in this section. $\alpha$ denotes the importance of the Dirichlet distribution.
(a) FedProx
(b) SCAFFOLD
Figure 9: Test accuracy and test loss on the CIFAR10 dataset with varying training temperatures with advanced federated learning schemes. Overall, models trained with lower temperatures showed superior performance.
Appendix D Experiment on hyperparameters
In this section, we explore the effect of other federated learning hyperparameters (i.e., batch size, number of local epochs, number of participants, and degree of data disparity). Through this set of experiments, we observed that applying low temperatures during local training improves test accuracy and convergence speed, being independent from hyperparameters. We note that the experiments were done by applying the default setting except for the control factors.
D.1 Batch size
(a) Batch size 8
(b) Batch size 16
(c) Batch size 32
(d) Batch size 64
(e) Batch size 128
Figure 10: Test accuracy and test loss over federated learning round with varying the batch size and the training temperature on the CIFAR10 dataset. Relatively small batch size shows better training quality with stability.
D.2 Number of participants
(a) Number of participants 5
(b) Number of participants 10
(c) Number of participants 15
Figure 11: Test accuracy and test loss over federated learning round with varying the number of participants and the training temperature on the CIFAR10 dataset. When relatively many clients participate in the training, it shows better training quality with stability
D.3 Number local epochs
(a) Number of local epoch 1
(b) Number of local epoch 5
(c) Number of local epoch 10
(d) Number of local epoch 15
Figure 12: Test accuracy and test loss over federated learning round with varying the number of local epochs and the training temperature on the CIFAR10 dataset. Relatively many updates in the local training phase show better training quality with stability
D.4 Degree of data disparity
(a) $\alpha$=0.1
(b) $\alpha$=0.5
(c) $\alpha$=1.0
Figure 13: Test accuracy and test loss over federated learning round with varying the degree of data disparity and the training temperature on the CIFAR10 dataset. Relatively moderate data disparity shows better training quality with stability
Hyperparameters Appendix D total clients 50 communication rounds 100 optimizer SGD learning rate 0.001 local epoch 5 participants per round 5 batch size 32 $\alpha$ 0.5 loss function Cross Entropy
Table 7: Default hyperparameters used in this section. $\alpha$ denotes the importance of the Dirichlet distribution.
Appendix E Details of Centered Kernel Alignment (CKA)
E.1 Background
Centered kernel alignment (CKA) is a method to compute the similarity between two different neural networks by measuring the similarity of the network representation for the given input data. Let $X$ and $Y$ be the features generated from two different neural networks, we can calculate the CKA by Eq 3. As we are assuming a federated learning scenario, the clients share the same neural network architecture. Thus, the dimensions of the two features from each corresponding layer are identical, allowing for a direct comparison to be possible. We note that in our experiments, we exploited the linear CKA to measure the similarity of the feature spaces between two neural networks.
$\begin{split}CKA(X,Y)&=\frac{||Y^{T}X||^{2}_{F}}{||X^{T}X||_{F}||Y^{T}Y||_{F}}\\ \end{split}$ (3)
E.2 Limitation of Centered Kernel Alignment evaluation in federated learning
Here, we discuss our rationale for measuring CKA between a client model participating in the federated learning round and a reference model trained using i.i.d. configurations. As mentioned in Section 4.4, the feature space similarity itself only reports the similarity of neural networks and does not assure high quality (i.e., accuracy) of the trained models. Figure 14 illustrates the possible cases that can observed in a federated learning scenario. We use these four cases to understand how locally trained models converge to the reference model. Here, the orange line indicates the update directions of local models from the federated learning training phase (i.e., diverging from the initially shared aggregated model presented as the black dot) and the red dotted line represents the average of these locally updated models that occur at the server. The star represents the ideal model in the representation space with high accuracy.
(a) Case 1
(b) Case 2
(c) Case 3
(d) Case 4
(e) Case 5
Figure 14: Illustration of possible cases in a federated learning scenario. The solid orange line and dotted red line correspond to the local update direction and aggregated model update direction respectively. The black circle and the red star depict the initial point and optimal point respectively.
$\bullet$ High similarity between local models does not guarantee a better model (Case 1, 2): The first case arises when local models, each trained on their respective local data, produce highly similar updates due to a well-aligned data configuration. Such alignment implies that the training data across different clients are relatively similar in distribution and characteristics, allowing the aggregated global model to closely approach the ideal model achievable through centralized training. However, in non-i.i.d. environments, achieving this scenario is infrequent. In non-i.i.d. settings, the data distribution varies significantly across different local nodes. This variation can lead to local models learning disparate aspects of the data, resulting in updates that are less similar. Consequently, the aggregated global model may not perform as well as in an i.i.d. scenario, as it has to reconcile these diverse and potentially conflicting updates. On the other hand, Case 2 represents a critical scenario in federated learning where there is a high similarity among local model updates, but these updates are aligned in a suboptimal or incorrect direction. This situation can be considered one of the worst cases in federated learning. In addition, using these two cases, we demonstrate that the high similarity between local models does not ensure a superior update or quality of training.
$\bullet$ A superior model can be obtained without similar local models (Case 3): Case 3 illustrates a case where a superior model can be achieved even when local models exhibit relatively low similarity. In this scenario, each local model learns effective but diverse features from heterogeneous data configurations, and these diverse updates are properly aggregated at the central server. Unlike scenarios where local models produce highly similar updates, in this case, the updates from different local models show significant variation. This diversity reflects the different aspects of the data each local model has learned. This situation arises under conditions of data heterogeneity, where the data distributions across different clients are diverse. Each local model captures unique features relevant to its specific data subset. Despite the low similarity among local model updates, the central server effectively aggregates these diverse updates. The aggregation process leverages the complementary information from each local model, resulting in a robust and well-generalized global model. This case supports the hypothesis that a proper global model can be obtained even when local models do not exhibit high similarity and is a globally ideal case for federated learning scenarios.
$\bullet$ Relationship between model similarity and the magnitude of updates (Case 4, 5): It is important to note that the magnitude of updates can influence the perceived similarity of the network. If the update magnitude is small, the similarity of the network may be measured as high, whereas if the update magnitude is large, the similarity may be measured as relatively low. This example underscores the need to rethink the conventional emphasis on high similarity as an indicator of performance. It suggests that focusing solely on similarity metrics, while providing insights on how local training operations induce model changes, might overlook other critical factors that contribute to the effectiveness of the aggregated model. Therefore, a more nuanced approach to evaluating model updates and their impact on the accuracy performance is required, considering both the magnitude of updates and their alignment with the overall learning objectives.
Based on the aforementioned discussions, we introduce measuring similarity with a reference model as an alternative evaluation method in this work. Although we cannot ensure that the reference model trained in an i.i.d. data configuration perfectly represents the optimal update objective, we consider this method as an effective and reasonable way to evaluate the federated learning process. This approach allows for a more nuanced assessment of the updates by comparing them to a standard benchmark, helping to identify whether the aggregated model is progressing toward a semi-optimal solution despite the inherent data heterogeneity.
E.3 Experimental results
Table 8 presents the exact numbers for the accuracy and CKA values discussed in Section 4.4 and Figure 5. Figure 15 illustrates the data distribution applied in the experiments.
Temperature ($T$) Accuracy (%) Layer 1 CKA Layer 2 CKA Layer 3 CKA Layer 4 CKA 4.00 49.72 0.6539 0.5632 0.7149 0.7661 2.00 53.82 0.6792 0.6872 0.7417 0.7963 1.00 58.19 0.7419 0.6997 0.7644 0.8214 0.50 63.48 0.9322 0.7459 0.8181 0.8559 0.25 66.85 0.9569 0.8161 0.8418 0.8830 0.05 72.76 0.9390 0.9059 0.8898 0.9026 Reference Model ($T$=1.00, i.i.d.) 73.86 1.0000 1.0000 1.0000 1.0000
Table 8: Accuracy and CKA values of 2 layer CNN measured on the CIFAR10 dataset in Figure 5.
Figure 15: Visualization of the distribution of training data used in CKA evaluation. Best viewed in color.
(a) $T$=0.05
(b) $T$=0.25
(c) $T$=0.5
(d) $T$=1.0
(e) $T$=2.0
(f) $T$=4.0
Figure 16: Client-wise feature space similarity matrix measured in the first convolution layers (Layer 1) by using CKA and inference accuracy in different temperatures (0.05, 0.25, 0.5, 1.0, 2.0, 4.0) and federated rounds (1, 5, 50, 100). (INIT: Initial dispatched model in the given round, AGG: Aggregated global model, REF: Reference model, ACC: Accuracy)
(a) $T$=0.05
(b) $T$=0.25
(c) $T$=0.5
(d) $T$=1.0
(e) $T$=2.0
(f) $T$=4.0
Figure 17: Client-wise feature space similarity matrix measured in the second convolution layers (Layer 2) by using CKA and inference accuracy in different temperatures (0.05, 0.25, 0.5, 1.0, 2.0, 4.0) and federated rounds (1, 5, 50, 100). (INIT: Initial dispatched model in the given round, AGG: Aggregated global model, REF: Reference model, ACC: Accuracy)
(a) $T$=0.05
(b) $T$=0.25
(c) $T$=0.5
(d) $T$=1.0
(e) $T$=2.0
(f) $T$=4.0
Figure 18: Client-wise feature space similarity matrix measured in penultimate layers (Layer 3) by using CKA and inference accuracy in different temperatures (0.05, 0.25, 0.5, 1.0, 2.0, 4.0) and federated rounds (1, 5, 50, 100). (INIT: Initial dispatched model in the given round, AGG: Aggregated global model, REF: Reference model, ACC: Accuracy)
(a) $T$=0.05
(b) $T$=0.25
(c) $T$=0.5
(d) $T$=1.0
(e) $T$=2.0
(f) $T$=4.0
Figure 19: Client-wise feature space similarity matrix measured in classification layers (Layer 4) by using CKA and inference accuracy in different temperatures (0.05, 0.25, 0.5, 1.0, 2.0, 4.0) and federated rounds (1, 5, 50, 100). (INIT: Initial dispatched model in the given round, AGG: Aggregated global model, REF: Reference model, ACC: Accuracy)